The purchase price of a natural​ gas-fired commercial boiler​ (capacity X) was ​$175,000 eight years ago. Another boiler of the same basic​ design, except with capacity 1.41X, is currently being considered for purchase. If it is​ purchased, some optional features presently costing $26,0000 would be added for your application. If the cost index was 170 for this type of equipment when the capacity X boiler was purchased and is 224 ​now, and the applicable cost capacity factor is 0.8, what is your estimate of the purchase price for the new​ boiler?
Answer #1

Cost of old boiler = 175000*224/170 = 230588.23

Cost of new boiler = 230588.23*(1.41X/X)^0.8 = 303538

Total cost = 303538+26000 = 329538